OpenStudy (anonymous):

Which of the following is equal to square root of the cube root of 2?

OpenStudy (anonymous):

\[\sqrt{\sqrt[3]{2}}\] thats the equation

OpenStudy (anonymous):

@jboy_52 : Options please: Other wise i think the answer is : \[\sqrt[3]{\sqrt{2}} = \sqrt[6]{2}\] 6th root of 2

OpenStudy (anonymous):

2 1/6 would be the answer?

OpenStudy (anonymous):

yes : 2 ^ 1/6 is a right answer too it's equal to: \[\sqrt[3]{\sqrt{2}} = \sqrt[3]{2^\frac{ 1 }{ 2}} = 2^\frac{ 1 }{ 3} \times 2^\frac{ 1 }{ 3} = 2^\frac{ 1 }{ 6}\]

OpenStudy (anonymous):

@jboy_52: Understood man..All concepts cleared?

OpenStudy (anonymous):

yea thank you man